OpenStudy (anonymous):

zayn malik scored 92 points. goals and assists are each worth a point. he scored 1/2 the number of assists than goals .what were his totals?

OpenStudy (anonymous):

let a=# assists, g=#goals. So, 92=g+a 92=g+(g/2) 184=2g+g 184=3g 61.33333=g a=g/2 a=61.33/2 a=30.66 These numbers don't make sense though because you can't score a fraction of a point!

OpenStudy (anonymous):

my teacher was having a bad day

OpenStudy (anonymous):

yeah, apparently :)